# Causal Feedback Discovery using Convergence Cross Mapping on Sea Ice Data

### v3

@@ 4.2.4. Convergent Cross Mapping (CCM)
CCM consistently identified the designed bidirectional feedbacks. Figures 58 illustrate results for $S_{2}\leftrightarrow S_{5}$: shadow manifolds (Figure 6), estimated cross-mapping (Figure 7), and convergence plots (Figure 8). Unlike Granger and PCMCI, CCM recovered weak causal links with high specificity.
Figure 5. Synthetic time series of $S_{2}$ and $S_{5}$.
[…]
Figure 6. Shadow manifolds $M_{x}$ and $M_{y}$ for different $\tau$.
[…]
Figure 7. Cross mapping performance ($S_{2}$ vs. $S_{5}$).
[…]
Figure 8. Convergence of CCM for $S_{1}\Leftrightarrow S_{5}$.
[…]
@@ 4.3. Real-World Arctic Climate Experiments
Next, we applied the abovementioned time series causal discovery methods to our Arctic climate dataset as shown in Table 1).
@@ 4.3.1. Granger Causality
Figure 9. Granger causality results on Arctic dataset.
[…]
@@ 4.3.2. PCMCI
Figure 10. PCMCI causal graph on Arctic dataset.
[…]
@@ 4.3.3. VarLiNGAM
VarLiNGAM identified robust linear causal relations (Table 3), including links between sea ice extent, wind, and humidity. While statistically strong, its scope remained restricted to linear effects.
Table 3. Causal Relationships identified by VarLiNGAM
Variable 1 Variable 2 Significance
$t2m(t)$ $\leftarrow$ sst(t1)sst(t-1) $(b>0)$ (100.0%)
specific_humidity(t)specific\_humidity(t) $\leftarrow$ wind_10m(t1)wind\_10m(t-1) $(b>0)$ (100.0%)
wind_10m(t)wind\_10m(t) $\leftarrow$ sea_ice_extent(t)sea\_ice\_extent(t) $(b>0)$ (100.0%)
$sst(t)$ $\leftarrow$ snowfall(t1)snowfall(t-1) $(b>0)$ (100.0%)
LW_down(t)LW\_down(t) $\leftarrow$ specific_humidity(t1)specific\_humidity(t-1) $(b>0)$ (100.0%)
$snowfall(t)$ $\leftarrow$ rainfall(t1)rainfall(t-1) $(b>0)$ (100.0%)
surface_pressure(t)surface\_pressure(t) $\leftarrow$ t2m(t1)t2m(t-1) $(b>0)$ (100.0%)
SW_down(t)SW\_down(t) $\leftarrow$ LW_down(t1)LW\_down(t-1) $(b>0)$ (98.0%)
@@ 4.3.4. CCM
Convergent Cross Mapping (CCM) revealed robust and physically consistent feedbacks that remain undetected by conventional linear approaches. In particular, CCM identified bidirectional couplings between sea ice extent and key climate variables, including snowfall, humidity, longwave radiation, and near-surface temperature. The detection of the snowfall–sea ice extent feedback, illustrated in Figures 1113, highlights the ability of CCM to uncover nonlinear interactions of critical relevance to Arctic climate dynamics.
Figure 11. Cross mapping: snowfall $\sim$ sea ice extent.
[…]
@@
Figure 12. Estimation: snowfall $\sim$ sea ice extent.
[…]
Cross mapping is evaluated using correlation, under the principle that higher correlation corresponds to stronger causal influence. As shown in the left panel of Figure 12, the performance of cross mapping from $M_{y}$ to $M_{x}$ for predicting $X(t)$ highlights that $X$ can be reliably inferred from $Y$. This indicates that $Y$ encapsulates sufficient dynamical information about $X$. Consequently, we infer that $X$ exerts a causal influence on $Y$. The key intuition is that if $X$ drives $Y$, then the historical dynamics of $Y$ inevitably encode traces of $X$, enabling robust causal inference through state-space reconstruction. For completeness, we report cross mapping in both directions; however, one direction consistently exhibits stronger predictive power, providing evidence of asymmetric causal dependence.
@@
Figure 13. Convergence: snowfall $\sim$ sea ice extent.
[…]
@@ 4.4. Comparative Performance Analysis
Tables 4 and 5 compare CCM and Granger using $p$-values.
For the sea ice dataset (Table 4), CCM detected significant feedbacks ($p=0.009$) for humidity, LW_down, SST, and t2m, consistent with known Arctic amplification processes. Granger, in contrast, reported universal significance ($p=0.000$), indicating false positives.
@@ 5. Discussion
While the findings validate theoretical understanding of Arctic climate processes, limitations remain. The synthetic dataset cannot fully represent the multiscale complexity of real climate systems, and this study focused only on the Arctic region. Future work should extend to diverse climatic contexts, explore hybrid approaches combining method strengths, and incorporate physical priors and uncertainty quantification to improve reliability of causal discovery in Earth system science.
@@ 6. Conclusions
This study provides the first systematic benchmarking of causal discovery methods on both synthetic and Arctic climate datasets. By comparing Granger causality, PCMCI, VarLiNGAM, and CCM, we establish that method performance is highly dependent on system dynamics. Granger and VarLiNGAM are effective for detecting strong linear dependencies, PCMCI improves robustness through conditional independence testing, but CCM emerges as the most valuable tool for complex climate systems due to its ability to capture nonlinear, coupled, and bidirectional feedbacks. Applied to Arctic data, CCM identified causal links between sea ice extent and key atmospheric variables (specific humidity, longwave radiation, and near-surface temperature), providing quantitative evidence for fundamental feedbacks driving Arctic amplification. These results underscore the importance of nonlinear methods in climate causal discovery and demonstrate the potential of CCM for advancing understanding of climate system interactions.
In the future, we would like to focus on extending CCM to high-dimensional settings, developing hybrid frameworks that integrate linear and nonlinear approaches, and applying these methods to other regions and processes critical to global climate dynamics. Continued refinement of causal discovery tools will improve the scientific basis for predicting and responding to climate change.

### v5

@@ 4.1.4. Convergent Cross Mapping (CCM)
CCM consistently identified the designed bidirectional feedbacks. Figures 6–Figures 9 illustrate results for $S_{2}\leftrightarrow S_{5}$: shadow manifolds (Figure 7), estimated cross-mapping (Figure 8), and convergence plots (Figure 9). Unlike Granger and PCMCI, CCM recovered weak causal links with high specificity.
Figure 6. Learned feedback by CCM between $S_{2}$ and $S_{5}$.:
Figure 7. Shadow manifolds $M_{x}$ and $M_{y}$ for different $\tau$.:
The shadow manifolds illustrated by Figure 7 shows the cross mapping of points from one manifold (blue box) to the corresponding points on another manifold (red star). If values from one manifold, for example $M_{X}$, produce a widely spread-out cross mapping to values on another manifold $M_{Y}$, this indicates that we are unable to precisely predict $M_{Y}$ (or the corresponding variable $Y$) given $M_{X}$. In such a case, limited information from $Y$ is embedded in $X$, suggesting that while $Y$ may drive $X$, the causal influence is weak. Conversely, the stronger the causality between variables, the narrower the cross mapping becomes. For instance, cross mapping $M_{X}$ from $M_{Y}$ (as shown in the left panel) results in a relatively narrow mapping, implying that $X$ has a stronger influence on $Y$. While a narrow cross convergence pattern can be an indicator of potential causality, it represents only one of the two key criteria required to establish causal relationships in Convergent Cross Mapping (CCM). The second criterion—convergence—requires that the accuracy of cross mapping improves as the library size increases.
Figure 8. Cross mapping performance ($S_{2}$ vs. $S_{5}$).:
We measure cross mapping using the correlation coefficient between the observed and predicted values. A stronger correlation indicates a stronger causal relationship. In Figure 8 the left panel illustrates the performance of cross mapping from manifold $M_{Y}$ to manifold $M_{X}$ in predicting $X(t)$. There exist little correlation. This means that we cannot accurately estimate $X$ given observations of $Y$, implying that $Y$ contains insufficient information about the dynamics of $X$. In such a case, we infer that $X$ weakly influences or causes $Y$.
Figure 9. Convergence of CCM for $S_{2}\Leftrightarrow S_{5}$.:
Figure 10. Learned causal graph with CCM on synthetic data:
Convergence means we are able to improve cross mapping accuracy the longer the period $L$ we consider. This happens only when we are able to enhance our reconstruction of the shared attractor between two variables using more data. In Figure 9, we find that both cross mappings converge, however one converges more than the other. We have satisfied the two criteria for CCM: convergence and cross mapping. We can thus say weak causality exists. However, the magnitudes of causality going from $S5$ or $S2$ are different.
@@ 4.2. Real-World Arctic Climate Experiments
Next, we applied the abovementioned time series causal discovery methods to our Arctic climate dataset as shown in Table 1.
@@ 4.2.1. Granger Causality
Figure 11. Granger causality results on Arctic dataset.:
@@ 4.2.2. PCMCI
Figure 12. PCMCI causal graph on Arctic dataset.:
@@ 4.2.3. VarLiNGAM
VarLiNGAM identified robust linear causal relations including links between sea ice extent, wind, and humidity. While statistically strong, its scope remained restricted to linear effects. Figure 13 is the learned causal graph from real-world data. We considered a lower threshold of $0.7$ meaning we considered edges with a causal effect larger than minimum threshold.
Figure 13. Learned causal graph from Arctic data using VarLiNGAM.:
@@ 4.2.4. CCM
Figure 14 revealed robust and physically consistent feedbacks present in the Arctic sea ice data that remain undetected by conventional linear approaches.
Figure 14. CCM learned causal graph on Arctic dataset.:
In particular, CCM identified bidirectional couplings between sea ice extent and key climate variables, including snowfall, humidity, longwave radiation, and near-surface temperature. The detection of the snowfall–sea ice extent feedback, illustrated in Figures 1517, highlights the ability of CCM to uncover nonlinear interactions of critical relevance to Arctic climate dynamics.
Figure 15. Cross mapping: snowfall $\sim$ sea ice extent.:
@@
Figure 16. Visualize correlation between snowfall $\sim$ sea ice extent.:
Physically, heavy snow on sea ice acts as an insulating layer that reduces heat loss from the ocean to the atmosphere, thereby slowing ice growth and reducing extent (a negative forcing on ice growth). Conversely, changes in sea ice extent can modify surface albedo, temperature, and moisture transport, thus influencing snowfall patterns (a feedback effect) (Clemens-Sewall et al., 2023). This is further seen in Figure 16 when CCM is applied. In the left panel the performance of cross mapping from $M_{y}$ to $M_{x}$ for predicting $X(t)$ highlights that $X$ can be reliably inferred from $Y$. This indicates that $Y$ encapsulates sufficient dynamical information about $X$. Consequently, we infer that $X$ exerts a causal influence on $Y$. The key intuition is that if $X$ drives $Y$, then the historical dynamics of $Y$ inevitably encode traces of $X$, enabling robust causal inference through state-space reconstruction. For completeness, we report cross mapping in both directions; however, one direction consistently exhibits stronger predictive power, providing evidence of asymmetric causal dependence.
@@
Figure 17. Convergence: snowfall $\sim$ sea ice extent.:
@@ 4.3. Comparative Performance Analysis
Tables 2 and 3 compare CCM and Granger using $p$-values. For the sea ice dataset (Table 2), CCM detected significant feedbacks ($p=0.009$) for humidity, LW_down, SST, and t2m, consistent with known Arctic amplification processes. Granger, in contrast, reported universal significance ($p=0.000$), indicating false positives.
@@ 5. Discussion
While the findings validate theoretical understanding of Arctic climate processes, limitations remain. The synthetic dataset cannot fully represent the multiscale complexity of real climate systems, and this study focused only on the Arctic region.
@@ 6. Conclusions
This study benchmarks causal discovery methods on synthetic and Arctic climate datasets. It compares Granger causality, PCMCI, VarLiNGAM, and CCM. Performance varies strongly with underlying system dynamics. Granger excels when dependencies are linear and strong. VarLiNGAM is also effective for linear, non-Gaussian structures. PCMCI improves robustness via conditional independence testing. CCM stands out in complex, nonlinear, and coupled systems. It can capture bidirectional feedbacks common in climate processes. Applied to Arctic observations, CCM revealed causal links involving sea ice extent. Specific humidity emerged as a key atmospheric driver. Longwave radiation showed strong causal interactions with sea ice. Near-surface temperature also exhibited influential feedback pathways. These findings provide quantitative evidence for mechanisms behind Arctic amplification. They highlight the necessity of nonlinear methods in climate causal discovery. Future work will extend CCM to high-dimensional settings, integrate hybrid linear–nonlinear frameworks, and apply them to other climate regions and processes.